OpenStudy (anonymous):

lim[x:1,(x^(3) - 1)/(x^(2)-6x+5)]

OpenStudy (asnaseer):

you the fact that you can factorise this expression as follows:\[\frac{x^3-1}{x^2-6x+5}=\frac{\cancel{(x-1)}(x^2+x+1)}{\cancel{(x-1)}(x-5)}=\frac{x^2+x+1}{x-5}\]

OpenStudy (anonymous):

thank so much....i made a mistake because (x-1) ( x^2 +2x+1)....:) god bless u

OpenStudy (asnaseer):

you're welcome

OpenStudy (anonymous):

thank u for ur help

OpenStudy (asnaseer):

no problem at all - I'm glad I was able to help :)

OpenStudy (anonymous):

:)